# SWIFT: A Monotonic, Flux-Form Semi-Lagrangian Tracer Transport Scheme for Flow with Large Courant Numbers

### v1

@@ 3.2 SWIFT Splitting for Consistent Tracers
First, the departure points and dry mass fluxes are recomputed using $\rho^{n}$ and the dry mass fluxes in equations (24)-(26) for the first transport step. The inner transport step for the tracer is then given by:
$(\rho m)^{x}$ $=(\rho m)^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},F^{x},\rho^{n}),$ (38a)
$(\rho m)^{y}$ $=(\rho m)^{n}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(m^{n},F^{y},\rho^{n}).$ (38b)
@@
$\begin{split}(\rho m)^{n+1}=(\rho m)^{n}&-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},F^{x},\rho^{n})-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{y},F^{x},\rho^{y})\\ &-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{n},F^{y},\rho^{n})-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{x},F^{y},\rho^{x}).\end{split}$ (40)
@@
$(\rho m)^{n+1}=\frac{1}{2}\left[(\rho m)^{y}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}(m^{y},F^{x},\rho^{y})+(\rho m)^{x}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}(m^{x},F^{y},\rho^{x})\right].$ (41)
@@ 3.3 SWIFT Splitting for Advective Tracers
$\begin{split}(\sigma m)^{n+1}=m^{n}&-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},u^{x})-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{y}_{A},u^{x},\sigma^{y})\\ &-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{n},u^{y})-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{x}_{A},u^{y},\sigma^{x}),\end{split}$ (43)
@@
$m^{x}_{A}$ $=(m^{n}-\Delta{t}\nabla_{x}\cdot F^{x}(m^{n},u^{x}))/\sigma^{x},$ (44)
$m^{y}_{A}$ $=(m^{n}-\Delta{t}\nabla_{y}\cdot F^{y}(m^{n},u^{y}))/\sigma^{y}.$ (45)
@@
$\sigma^{n+1}=\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x})-\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y}).$ (47)
@@ 3.4.1 Consistency
$(\rho m)^{x}$ $=\rho^{n}K-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(K,F^{x},\rho^{n})=\rho^{n}K-\Delta{t}K\nabla_{x}\cdot F^{x}=K\rho^{x},$ (49a)
$(\rho m)^{y}$ $=\rho^{n}K-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(K,F^{y},\rho^{n})=\rho^{n}K-\Delta{t}K\nabla_{y}\cdot F^{y}=K\rho^{y},$ (49b)
@@ 3.4.2 Multi-dimensional Monotonicity Properties of One-Dimensional Schemes
$(\rho m)^{n+1}=(\rho m)^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},F^{x},\rho^{n}),$ (52)
@@
$(\rho m)^{n+1}=\frac{1}{2}(\rho m)^{x}-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{x},F^{y},\rho^{x})+\frac{1}{2}(\rho m)^{y}-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{y},F^{x},\rho^{y}).$ (53)
@@ 3.4.3 Stability and Relationship to COSMIC
$\rho^{n+1}=\rho^{n}-\frac{\Delta t}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x})-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{y}_{A},u^{x})-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y})-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{x}_{A},u^{y}).$ (55)
Comparing (55) with (35) reveals that, in the absence of a limiter, the COSMIC and SWIFT splittings differ through the inclusion of $\sigma^{x}$ and $\sigma^{y}$ arguments for the outer flux evaluations and thus the computation of the departure points for the outer step.
A corollary of this is that if the transporting velocity field is constant in space, then the COSMIC and SWIFT splittings in Sections 2.2.1 and 3.1 for density are equivalent. This can be seen by noting that $\sigma^{x}=\sigma^{y}=1$ if the flow is constant.
@@ 4 Two-Dimensional Results
We test the SWIFT splitting by running a number of prescribed velocity transport tests on the doubly-periodic plane and comparing with COSMIC splitting. For each test we use PPM as the one-dimensional scheme, and where noted use the strict limiting. For each test the density $\rho$ is computed using the same splitting as the tracer $m$. All tracers are transported conservatively, have an initial minimum of 0 and maximum of 1. The set up of the test cases is given in Appendix B. Throughout the results it is demonstrated that at small Courant numbers, the behaviour of the SWIFT and COSMIC splittings are very similar and so SWIFT retains the benefits of the COSMIC splitting. However at large Courant numbers, the COSMIC splitting does not maintain monotonicity when a limiter is used, while the SWIFT splitting does.
@@ 4.1 Constant Velocity Test
Table 1: Results for the constant velocity tests. The top set are for a constant density field, the bottom are for the varying density. The normalised $L^{2}$ errors, minimum and maximum values are given for slotted cylinder initial conditions on the $128\times 128$ grid, for Courant numbers $c_{max}=0.256$ and $c_{max}=2.56$. The mixing ratio $m$ has no limiter applied, while $m^{L}$ uses the limiter of Section A.2.
Statistics from the end of these simulations (at $t=1000s$) are shown in Table 1. The “Min” and “Max” columns give the instantaneous minima and maxima of the transported fields at the end of the final time step. The tracer field has an initial minimum of 0 and an initial maximum of 1, while the spatially-varying density has an initial minimum of 0.6 and an initial maximum of 1. The “$L^{2}$ error” column describes the normalised $L^{2}$ error, which is given by $||q-q_{true}||/||q_{true}||$ for transported field $q$ and analytic solution $q_{true}$. The left-hand side of the table presents the values from a run with the smaller Courant number, while the right-hand side contains the values for the larger Courant number case. Each row represents a different transported variable. The superscript for $m^{L}$ indicates that the monotonic limiter was used for the transport of that mixing ratio. The upper half of the table contains statistics for tracers transported relative to an initially-constant density, while the lower half describes the case with a spatially-varying density. The final fields corresponding to the lower-right corner of the table are plotted in Figure 2.
@@ 4.2 Non-Divergent Deformational Test
Figure 3: The transported fields at the end of the final time step ($t=1000$ s) of the non-divergent deformational test of Section 4.2 with a constant density. These fields are taken from the large Courant number case, with $\Delta t=2$ s. Different rows show the fields transported with the COSMIC splitting and SWIFT splittings. The mixing ratio fields have contours every $0.1$ kg kg-1, omitting the zero contour. Arrows on the density plots show the direction and magnitude of the transporting velocity.
@@
Figure 4: The transported fields at the halfway point ($t=500$ s, top two rows) and at the end of the final time step ($t=1000$ s, bottom two rows) of the non-divergent deformational test of Section 4.2. These fields are taken from the large Courant number case, with $\Delta t=2$ s. Different rows show the fields transported with the COSMIC splitting and SWIFT splittings. Even under strong deformation, the limited tracer with the SWIFT splitting remains monotonic. The density fields are contoured at intervals of $0.05$ kg m-3, with the $0.8$ kg m-3 contour omitted. The mixing ratio fields have contours every $0.1$ kg kg-1, omitting the zero contour. Arrows on the density plots show the direction and magnitude of the transporting velocity.
@@
Table 2: Statistics from the non-divergent deformational test. The normalised $L^{2}$ errors, minimum and maximum values are given for the slotted cylinder initial conditions on the $128\times 128$ grid, for Courant numbers $c_{max}=0.512$ and $c_{max}=5.12$. The mixing ratio $m$ has no limiter applied, while $m^{L}$ uses the limiter of Section A.2.
@@ 4.3 Divergent Deformational Test
The non-divergent deformational flow can be modified to make a challenging divergent test case (see Appendix for test case set up). The initial density is a smooth field, while the initial tracers are again slotted cylinders. The divergent flow stretches out the tracers, before returning them to their initial conditions. As with the previous two tests, this was run with two different time steps: $\Delta t=0.2$ s which corresponds to $c_{max}=0.383$, and $\Delta t=2$ s which gives $c_{max}=3.83$.
Table 3 contains results from this test. The table is structured in the same way as Tables 1 and 2, and demonstrates that with divergent flow the SWIFT splitting still maintains monotonicity properties, while error norms are again comparable with those of the COSMIC splitting. Fields at $t=500$ s and $t=1000$ s from the large Courant number case are plotted in Figure 5.
Figure 5: The transported fields at the halfway point ($t=500$ s, top two rows) and end of the final time step ($t=1000$ s, bottom two rows) of the divergent test of Section 4.2. These fields are taken from the large Courant number case, with $\Delta t=2$ s. Different rows show the COSMIC and SWIFT splittings. The density fields are contoured at intervals of $0.05$ kg m-3, with the $0.8$ kg m-3 contour omitted. The mixing ratio fields have contours every $0.1$ kg kg-1, omitting the zero contour. Arrows on the density plots show the direction and magnitude of the transporting velocity.
@@
Table 3: Results for the divergent transport test. The normalised $L^{2}$ errors, final minimum and maximum values are given for the slotted cylinder initial conditions on the $128\times 128$ grid, for Courant numbers $c_{max}=0.383$ and $c_{max}=3.83$. The mixing ratio $m$ has no limiter applied, while $m^{L}$ uses the limiter of Section A.2.
@@ 4.4 Convergence Tests
We investigated the convergence of the schemes in different regimes. Using the constant velocity test of Section 4.1 and the non-divergent deformation test of 4.2, the relative $L^{2}$ error norm was measured at different spatial resolutions whilst holding the Courant number constant (by also varying the time step). This was done for two different maximum Courant numbers ($0.256$ and $2.56$ for test 1, $0.6$ and $6$ for test 2) and on grids with $64^{2}$, $128^{2}$, $256^{2}$ and $512^{2}$ cells. Convergence rates for different variables are shown in Table 4. All convergence rates are determined by fitting a line of best fit between the logarithm of the grid spacing and the relative $L^{2}$ error. For the constant velocity (test 1), the convergence rates for the density field and for the tracer field transported relative to a constant density achieve the theoretical third-order values (for the unlimited case). However when the density is spatially-varying, the tracer field has a second-order convergence rate for both the COSMIC and SWIFT splittings. We believe that this is because the departure distance calculation of (26) corresponds to an assumption that the density is constant within a cell. Future work will investigate the impact of using a higher-order subgrid reconstruction of the density to compute the departure distance. For the deformational velocity (test 2) with large maximum Courant number, the convergence rates are all approximately second-order. This is because at large Courant numbers with non-constant velocity, the second-order splitting error of the COSMIC and SWIFT splitting dominates.

### v2

@@ 3.2 SWIFT Splitting for Consistent Tracers
First, the departure points and dry mass fluxes are recomputed using $\rho^{n}V$ and the dry mass fluxes substituted into equations (9)-(11) for the first transport step. The inner transport step for the tracer is then given by:
$(\rho m)^{x}$ $=(\rho m)^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},F^{x},\rho^{n}V),$ (41a)
$(\rho m)^{y}$ $=(\rho m)^{n}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(m^{n},F^{y},\rho^{n}V).$ (41b)
@@
$\begin{split}(\rho m)^{n+1}=(\rho m)^{n}&-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},F^{x},\rho^{n}V)-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{y},F^{x},\rho^{y}V)\\ &-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{n},F^{y},\rho^{n}V)-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{x},F^{y},\rho^{x}V).\end{split}$ (43)
@@
$(\rho m)^{n+1}=\frac{1}{2}\left[(\rho m)^{y}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}(m^{y},F^{x},\rho^{y},V)+(\rho m)^{x}-\Delta t\nabla_{y}\cdot\mathcal{F}^{y}(m^{x},F^{y},\rho^{x},V)\right].$ (44)
@@ 3.3 SWIFT Splitting for Advective Tracers
$\begin{split}(\sigma m)^{n+1}=m^{n}&-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},u^{x},V)-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{y}_{A},u^{x},\sigma^{y}V)\\ &-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{n},u^{y},V)-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{x}_{A},u^{y},\sigma^{x}V),\end{split}$ (46)
@@
$m^{x}_{A}$ $=(m^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},u^{x},V))/\sigma^{x},$ (47)
$m^{y}_{A}$ $=(m^{n}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(m^{n},u^{y},V))/\sigma^{y}.$ (48)
@@
$\sigma^{n+1}=\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x},V)-\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y},V).$ (50)
@@ 3.4 Analysis of SWIFT Splitting
3.4.1 Divergence-Linearity
The SWIFT splitting for density transport presented in Section 3.1 provides increments that are linear in the divergence. This can be seen by inserting $\rho^{n}=K$ for constant $K$ into (35), which when combined with (36) returns densities for the outer step of $\rho_{A}^{x}=K$ and $\rho_{A}^{y}=K$. These values can then be substituted into (38). Using the property discussed in Section 2.1.3 that when the density is constant, the mass flux is proportional to the wind field, the resulting density is
$\rho^{n+1}=K(1-\Delta t\bm{\nabla\cdot u}).$ (51)
Thus, the increment to the field is linear in the divergence, and a constant density is preserved under divergence-free flow.
@@ 3.4.2 Consistency
$(\rho m)^{x}$ $=\rho^{n}K-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(K,F^{x},\rho^{n}V)=\rho^{n}K-\Delta{t}K\nabla_{x}\cdot F^{x}=K\rho^{x},$ (53a)
$(\rho m)^{y}$ $=\rho^{n}K-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(K,F^{y},\rho^{n}V)=\rho^{n}K-\Delta{t}K\nabla_{y}\cdot F^{y}=K\rho^{y},$ (53b)
@@ 3.4.3 Multi-dimensional Monotonicity Properties of One-Dimensional Schemes
$(\rho m)^{n+1}=(\rho m)^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},F^{x},\rho^{n}V),$ (56)
@@
$(\rho m)^{n+1}=\frac{1}{2}(\rho m)^{x}-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(m^{x},F^{y},\rho^{x}V)+\frac{1}{2}(\rho m)^{y}-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(m^{y},F^{x},\rho^{y}V).$ (57)
@@ 3.4.4 Stability and Relationship to COSMIC
$\begin{split}\rho^{n+1}=\rho^{n}&-\frac{\Delta t}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x},V)-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{y}_{A},u^{x},V)\\ &-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y},V)-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{x}_{A},u^{y},V).\end{split}$ (59)
This can be compared directly with the transport of density using SWIFT, and so repeating (39) here
$\begin{split}\rho^{n+1}\equiv\rho^{n}&-\frac{\Delta t}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x},V)-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y}V)\\ &-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y},V)-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x}V),\end{split}$ (60)
reveals that, in the absence of a limiter, the COSMIC and SWIFT splittings differ through the inclusion of $\sigma^{x}$ and $\sigma^{y}$ arguments for the outer flux evaluations and thus the computation of the departure points for the outer step.
A corollary of this is that if the transporting velocity field is constant in space, then the COSMIC splitting of Section 2.3.1 and the SWIFT splitting of Section 3.1 for density are equivalent. This can be seen by noting that $\sigma^{x}=\sigma^{y}=1$ if the flow is constant.
@@ 4 Two-Dimensional Results
We test the SWIFT splitting by running a number of prescribed velocity transport tests on the doubly-periodic plane with $x$ and $y$ ranging from $-0.5$ km to $0.5$ km, and comparing with COSMIC splitting. For each test we use PPM as the one-dimensional scheme, and where noted use the strict limiting. For each test the density $\rho$ is computed using the same splitting as the tracer $m$. All tracers are transported conservatively, have an initial minimum of 0 and maximum of 1. The setup of the test cases is given in Appendix B, but in all tests the scalar fields are transported once around the domain so that the analytic final solutions are equal to their initial values. Throughout the results it is demonstrated that at small Courant numbers, the behaviour of the SWIFT and COSMIC splittings are very similar and so SWIFT retains the benefits of the COSMIC splitting. However at large Courant numbers, the COSMIC splitting does not maintain monotonicity when a limiter is used, while the SWIFT splitting does.
@@ 4.1 Constant Velocity Test
Table 1: Results for the constant velocity tests. The top set are for a constant density field, the bottom are for the varying density. The normalised $L^{2}$ errors, minimum and maximum values are given for slotted cylinder initial conditions on the $128\times 128$ grid, for Courant numbers $c_{max}=0.256$ and $c_{max}=2.56$. The mixing ratio $m$ has no limiter applied, while $m^{L}$ uses the limiter of Appendix AA.2.
Statistics from the end of these simulations (at $t=100$ s) are shown in Table 1. The “Min” and “Max” columns give the instantaneous minima and maxima of the transported fields at the end of the final time step. The tracer field has an initial minimum of 0 and an initial maximum of 1, while the spatially-varying density has an initial minimum of 0.6 and an initial maximum of 1. The “$L^{2}$ error” column describes the normalised $L^{2}$ error, which is given by $||q-q_{true}||/||q_{true}||$ for transported field $q$ and analytic solution $q_{true}$. The left-hand side of the table presents the values from a run with the smaller Courant number, while the right-hand side contains the values for the larger Courant number case. Each row represents a different transported variable. The superscript for $m^{L}$ indicates that the monotonic limiter was used for the transport of that mixing ratio. The upper half of the table contains statistics for tracers transported relative to an initially-constant density, while the lower half describes the case with a spatially-varying density. The final fields corresponding to the lower-right corner of the table are plotted in Figure 2.
@@ 4.2 Non-Divergent Deformational Test
Figure 3: The transported fields at the end of the final time step ($t=100$ s) of the non-divergent deformational test of Section 4.2 with a constant density. These fields are taken from the large Courant number case, with $\Delta t=2$ s. Different rows show the fields transported with the COSMIC splitting and SWIFT splittings. The mixing ratio fields have contours every $0.1$ kg kg-1, omitting the zero contour. Arrows on the density plots show the direction and magnitude of the transporting velocity.
@@
Figure 4: The transported fields at the halfway point ($t=50$ s, top two rows) and at the end of the final time step ($t=100$ s, bottom two rows) of the non-divergent deformational test of Section 4.2. These fields are taken from the large Courant number case, with $\Delta t=2$ s. Different rows show the fields transported with the COSMIC splitting and SWIFT splittings. Even under strong deformation, the limited tracer with the SWIFT splitting remains monotonic. The density fields are contoured at intervals of $0.05$ kg m-3, with the $0.8$ kg m-3 contour omitted. The mixing ratio fields have contours every $0.1$ kg kg-1, omitting the zero contour. Arrows on the density plots show the direction and magnitude of the transporting velocity.
@@
Table 2: Statistics from the non-divergent deformational test. The normalised $L^{2}$ errors, minimum and maximum values are given for the slotted cylinder initial conditions on the $128\times 128$ grid, for Courant numbers $c_{max}=0.512$ and $c_{max}=5.12$. The mixing ratio $m$ has no limiter applied, while $m^{L}$ uses the limiter of Appendix AA.2.
@@ 4.3 Divergent Deformational Test
The non-divergent deformational flow can be modified to make a challenging divergent test case (see Appendix B for the test case setup). The initial density is a smooth field, while the initial tracers are again slotted cylinders. The divergent flow stretches out the tracers, before returning them to their initial conditions. As with the previous two tests, this was run with two different time steps: $\Delta t=0.2$ s which corresponds to $c_{max}=0.383$, and $\Delta t=2$ s which gives $c_{max}=3.83$.
Table 3 contains results from this test. The table is structured in the same way as Tables 1 and 2, and demonstrates that with divergent flow the SWIFT splitting still maintains monotonicity properties, while error norms are again comparable with those of the COSMIC splitting. Fields at $t=50$ s and $t=100$ s from the large Courant number case are plotted in Figure 5.
Figure 5: The transported fields at the halfway point ($t=50$ s, top two rows) and end of the final time step ($t=100$ s, bottom two rows) of the divergent test of Section 4.2. These fields are taken from the large Courant number case, with $\Delta t=2$ s. Different rows show the COSMIC and SWIFT splittings. The density fields are contoured at intervals of $0.05$ kg m-3, with the $0.8$ kg m-3 contour omitted. The mixing ratio fields have contours every $0.1$ kg kg-1, omitting the zero contour. Arrows on the density plots show the direction and magnitude of the transporting velocity.
@@
Table 3: Results for the divergent transport test. The normalised $L^{2}$ errors, final minimum and maximum values are given for the slotted cylinder initial conditions on the $128\times 128$ grid, for Courant numbers $c_{max}=0.383$ and $c_{max}=3.83$. The mixing ratio $m$ has no limiter applied, while $m^{L}$ uses the limiter of Appendix AA.2.
@@ 4.4 Convergence Tests
We investigated the convergence of the schemes in different regimes. Using the constant velocity test of Section 4.1 and the non-divergent deformation test of 4.2, the relative $L^{2}$ error norm was measured at different spatial resolutions whilst holding the Courant number constant (by also varying the time step). This was done for two different maximum Courant numbers ($0.256$ and $2.56$ for test 1, $0.6$ and $6$ for test 2) and on grids with $64^{2}$, $128^{2}$, $256^{2}$ and $512^{2}$ cells. Convergence rates for different variables are shown in Table 4. All convergence rates are determined by fitting a line of best fit between the logarithm of the grid spacing and the relative $L^{2}$ error. For the constant velocity (test 1), the convergence rates for the density field and for the tracer field transported relative to a constant density achieve the theoretical third-order values (for the unlimited case). However when the density is spatially-varying, the tracer field has a second-order convergence rate for both the COSMIC and SWIFT splittings. We believe that this is because the departure displacement calculation of (11) for the transport of tracers corresponds to an assumption that the density is constant within a cell. Future work will investigate the impact of using a higher-order subgrid reconstruction of the density to compute the departure displacements. For the deformational velocity (test 2) with large maximum Courant number, the convergence rates are all approximately second-order. This is because at large Courant numbers with non-constant velocity, the second-order splitting error of the COSMIC and SWIFT splitting dominates.